OpenStudy (anonymous):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x). A. 7 + 4x B. 4 – x C. –x – 4 D. –4 – 7x

OpenStudy (anonymous):

\[f(x) + g(x)=3x-2+6-4x\] \[f(x)+g(x)=-x+4\]

OpenStudy (anonymous):

which is the same as writing \[f(x)+g(x)= 4-x\]

OpenStudy (anonymous):

do you understand what i have done??

OpenStudy (anonymous):

g(x)=6-4x f(x)=3x-2 f(x)+g(x)=(6-4x)+(3x-2) f(x)+g(x)=(4-x)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw....